OpenStudy (anonymous):

solve for p, show your work 5p + 6 = -4p - 8

OpenStudy (anonymous):

Take all terms containing p to one side, and all the numbers to the other, what do you get? 5p + 4p = -8 -6 Now solve it.

OpenStudy (anonymous):

the 4p loses its negative when it switches sides?

OpenStudy (anonymous):

Erm...sort of...you just add 4p to both sides so (-4p+4p)=0 and cancelled out...

OpenStudy (anonymous):

Just remember, when you add one thing to one side, you -must- add it to the other side to keep the equation balanced